# Illicit Promotion on Twitter

### v2

@@ 4. Posts of Illicit Promotion
The ratios of PIPs to the Twitter stream. Another interesting question is to profile what fraction of general tweets are PIPs. As described in §3.1, we utilized daily tweet snapshots from the Twitter Archiving Project to evaluate the generalizability of our binary PIP classifier, which achieves a precision of 99.33%. We then selected the daily tweet snapshot of the 6th day of each month between January 2023 and June 2023, and applied our binary PIP classifier to these 6 snapshots. As presented in Table 5, the ratio of PIPs ranges from 2.69% to 4.75%, which, as a concerning fraction, highlights the prevalence of PIPs as well as the necessity of more mitigation efforts.
Categories. Leveraging the multiclass PIP classifier (§3.2), all the captured PIPs have been predicted into one of well-defined 11 categories of illicit goods and services. The distribution of PIPs across these categories is listed in Table 6. And we can see the top 3 categories are Pornography,Gambling and Illegal Drug, which account for 91.17% PIPs in sum. Belong, we provide more observations regarding the categories of PIPs.
Table 6. The distribution of PIPs across illicit goods and services categories.
@@
Gambling 13.34% Forgery and Fake Documents 0.43%
@@
Data Theft and Leakage 2.03% Harassment 0.18%
@@
Among these categories, pornography is the most prevalent one, with 69.78% PIPs belonging to this category. Particularly, we found out that many PIPs were used to promote products or services of child pornography, which have severely violated Twitter’s content policy regarding child sexual exploitation (chi, 2023). Also, to evade detection, many pornography PIPs contain only innocent or seemingly harmless text. Instead, the images attached to these tweets contain necessary visual elements that both hint child pornography as well as providing explicit instructions regarding how to access child pornography. For instance, as shown in Figure3(a), the tweet text does not reveal itself as pornography, nor does it specify how to access the resources for child pornography, but the attached image states that the user needs to go to the author’s homepage to learn the telegram group(Figure 3(b)).
(a) A PIP with an image attached to signal child pornography.
(b) The account profile of the PIP presented in Figure 3(a).
Figure 3. An example of a pornography PIP.
Additionally, geographical names are often used in PIPs so as to advertise location-based illegal services, e.g., local sex services. For example, the following two tweets from two different accounts, “Candice Bridges Elton CopperField #Guangzhou #Guangzhou Massage” and “Lesley Browne Baldwin Yerkes #Guangzhou #Guangzhou Massage”, both use massage and wellness as a way to promote illegal sex services in Guangzhou, China. Also, both differ only in the text but share the same images, hashtags, geographical names, and contacts.
Besides, 8.04% PIPs involve the promotion of various drugs, e.g., methamphetamine, marijuana, and heroin. As illustrated in Figure 4, these tweets promote drugs in Chinese and Thai, respectively. Also, these PIPs have utilized jargon words. For example, ”猪肉”, a Chinese word denoting pork, is used in illicit promotion to name methamphetamine. It is indeed surprising to see them get posted successfully on Twitter since even the tweet text itself appears sufficient to decide it is drug-related, not to mention the hashtags as well as the images. Besides, weapon PIPs often specify the detailed model of the weapons on sale. And some even include images to demonstrate the item’s availability. Similarly, harassment PIPs are mainly used to promote harassment as a service, such as SMS bombing, and phone call bombing. Lastly, PIPs in the category of data theft&leakage encompass many products, such as hacking services, stolen account credentials, unauthorized access to confidential information, and personal data leakage. Concrete PIP examples for these categories can be found in Appendix A.
(a) A drug promotion for methamphetamine in Chinese.
(b) A drug promotion for marijuana in Thai.
Figure 4. Examples of drug PIPs.
@@
We further investigate the accounts that have all posts detected as PIPs, which comprise 7.87% of all PIP accounts. Our analysis reveals that such accounts tend to be dedicated to promoting a specific category of PIPs. Also, some accounts have posted very few tweets since registration, but rely primarily on their profiles rather than tweets for illicit promotion. Also, among top accounts with the most PIPs, we observe that a Twitter account can post hundreds of thousands of PIPs without getting banned by the platform. For instance, a PIP account registered in February, 2020 has posted over 198k PIPs in the category of data leakage, which strongly indicates the limitations of Twitter’s content moderation.
PIP engagements. Over time, PIPs accumulate engagements such as likes, replies, retweets and quotes (i.e. retweets with comments). Given the date a PIP is published $t_{p}$ and the date it is crawled $t_{c}$, $t_{e}=t_{c}-t_{p}$ is the elapse time of a PIP, i.e., the days passed since it is posted. As described in §5, 90% PIPs survive the first two months, thus we group PIPs based on elapse time ranging from 1 day to 60 days, and calculate the average engagement of each group. As shown in Figure 6, PIPs indeed receive a non-negligible volume of engagement, e.g., PIPs can receive 374 likes on average after being posted for 30 days, while the average number of likes a regular tweet can receive is 37 (ave, 2021). What is also observed is that the extent of engagement declines gradually after 35 days. We believe it is because PIPs with more engagement are more likely to be detected and unpublished by Twitter, which renders the survivorship bias wherein only the surviving cases will be counted for groups of long elapse time. However we are not clear about the reason why average retweets show significant fluctuations after 40 days.
Figure 6. Average engagements of PIPs over time.
@@
PIP distribution across natural languages. To recognize the natural language of a PIP, fastText  [^4] https://fasttext.cc/, a library for text representation and classification from Facebook’s AI Research (FAIR) lab, is applied. As the result, most PIPs (over 99%) are grouped into 5 major natural languages including Chinese, English, Japanese, Thai and Spanish. In terms of illicit categories, the top 5 languages are similar to each other as shown in Table 7. Pornography takes a major place in almost all of the top languages, followed by gambling or illegal drug. Considering products of these categories are illegal but can be of great demand in most countries, it’s reasonable to have observed such results.
Table 7. Top-5 languages of PIPs.
Languages % PIP Top3 Categories
Chinese 48.75% Pornography Gambling Illegal Drug
English 38.24% Pornography Gambling Illegal Drug
Japanese 9.98% Pornography Gambling Illegal Drug
Thai 1.70% Illegal Drug Surrogacy Pornography
Spanish 0.54% Illegal Drug Pornography Gambling
Finding I: Millions of illicit promotion posts have been observed on the Twitter platform, which are widely distributed in terms of Twitter accounts, categories and products of illicit goods and services, and natural languages.
Evasion tactics of PIPs. To comprehensively analyze the evasion tactics of PIPs, we randomly sample 500 PIPs for each illicit category, manually look into each of them to distll tactics, and implement quantitative analysis on the whole PIP dataset to verify the applicability. The tactics adopted by miscreants to evade PIP detection can be summarized into four categories, as detailed below.
Jargon words. Across PIPs of different categories, jargon words are commonly used. Particularly, through looking into PIPs, we have manually identified a set of 108 different jargon words that are embedded into 31.21% of all PIPs, which can only serve as a lower-bound estimate for the adoption of jargon words in PIPs. Also, the adoption of jargon words is observed for all the illicit categories. For example, in Chinese PIPs of drugs, ”叶子” (leaves in English), or its emoji is used to refer to marijuana, while ”猪肉” (pork) and ”冰”(ice) and their emojis are used to represent methamphetamine. These terms are derived from either the color or the shape of the respective illegal items. Additionally, a metaphor of the nature of the activity is also popular as jargon words for PIPs. For instance, the farmer and its emoji are used in drug PIPs to refer to people who plant marijuana. The use of jargon words helps illicit promotion operators blend their content with benign tweets, thus impeding content moderation. More jargon examples can be found in Appendix B.
Finding II: Illicit promotion campaigns have adopted various evasion tactics, likely in an attempt to evade content moderation of the Twitter platform.
@@
Availability across online social networks To investigate the availability of PIPs across social network platforms, we conducted a cross-platform analysis for other three major platforms beyond Twitter: YouTube, Facebook, and TikTok. For each PIP category, we randomly selected 20 PIP keywords which led to PIPs of the respective category identified on Twitter, and manually searched these platforms for PIPs if any. As shown in Table 8, PIPs are present on all these three platforms, albeit with varying degrees of prevalence. Particularly, Facebook has the highest hit ratio, with 83% of the sampled search keywords having at least one PIP located, and Youtube with the lowest at 35%. Besides, the availability of PIPs varies across platforms and categories. For example, the hit ratios of both gambling and data theft and leakage are higher on Tiktok than other subcategories. On the other hand, the promotion of weapon sales has a low hit ratio on all three platforms, with less than 30% of the sampled seeds having one or more PIPs observed. Interestingly, we found that some PIPs on different platforms share the same contact, indicating that they belong to the same promotion campaign.
Finding III: Illicit promotion has also been observed at a concerning prevalence for other three popular OSNs including Facebook, Youtube, and TikTok.
@@ 5. Content Moderation Against Illicit Promotion
All OSNs under our study claim to have enforced strict content moderation (tik, 2023; twi, 2023b; you, 2023), in which case, violative posts should either be blocked from publishing or get unpublished once detected in a later time. However, The existence of so many violative PIPs of diverse illicit goods and services categories suggest respective OSNs especially Twitter fail to prevent PIPs from being posted and becoming visible to OSN users. In this section, we present more in-depth observations regarding Twitter’s content moderation towards PIPs and PIP accounts.
Table 9. The evasion rates of PIPs across time.
@@
To further profile the effectiveness of Twitter’s content moderation towards published PIPs, we carried out revisiting for captured PIPs. Specifically, 50,000 PIPs hunted in each round are sampled, and their availability are tested by revisiting them periodically (usually at a weekly pace). Given sampled $PIP_{date_{p}}$ which are first posted on date $date_{p}$, their evasion rate $ER_{date_{p}\text{, }date_{r}}$ is defined as the ratio of PIPs that are still reachable when being revisited on date $date_{r}$. Table 9 presents the revisiting results between April 10 and May 1 in 2023, for PIPs posted between Oct 24, 2022 and March 12, 2023. And we can see PIPs posted between Oct 24-30, 2022 have 21.46% still reachable when revisiting 6 month later in the end of April 2023. Also, more than 90% PIPs can survive the first two months since being published.
@@
We then investigate why some PIPs become unavailable during revisits and find out that most are due to account suspension. Specifically, Twitter returns one of six error messages if the PIP is unavailable. These messages can tell us the underlying reasons of PIP unavailability. For instance, a PIP may be unpublished due to the suspension of the parental account, in which case, the error message will be ”This Tweet is from a suspended account.”. Other reasons include page non-existence, deletion by the author, account non-existence and violation of Twitter’s rules. Among PIPs unreachable during revisits, 91.59% are due to account suspension, and 6.22% are due to page non-existence. In summary, we can see that Twitter works in a continuous manner to detect and suspend PIP accounts, resulting in tweets (including PIPs) of the detected PIP accounts also become unpublished.
@@
However, many PIPs can still survive for a long period, and we further investigate the difference between unpublished PIPs and surviving ones. In an attempt to answer this question, PIPs are sampled and divided into two groups depending on the length of their evasion time as observed during revisiting. We then compare both groups with regards to various aspects, e.g., PIP categories, the text syntactic and semantics, the writing style, characteristics of the posting accounts. The only significant difference we have observed resides in the number of PIPs posted by the parental Twitter accounts. For the group of PIPs with a short lifetime, their parental Twitter accounts have 379 PIPs observed on average in our dataset. On the contrary, it is only 62 for the group of PIPs with much longer evasion time. A reasonable explanation is that the more PIPs a Twitter account has posted, the more likely it will be captured and thus suspended by the platform, in which case, all its tweets will also be unavailable, including the PIPs. To further profile this observation, we sampled 153,921 out of all the observed 580,530 PIP accounts, and revisited them during April 15-19, 2023. By then, only 78.95% accounts were still available while all the others got suspended. Comparing alive PIP accounts with suspended ones reveals that alive PIP accounts have 3 PIPs observed on average while blocked ones have 87. Finding IV: An arms race is observed between illicit promotion campaigns and Twitter content moderation, as the result of which, almost 80% PIPs have got banned six months later after being published while on the other hand, 90% PIPs can survive the first two months.
@@ 6.1. PIP Contacts
Evasion techniques for promoting PIP contacts. As discussed in §4, PIP operators have adopted various evasion or promotion tactics so as to make the resulting PIPs appear benign. Furthermore, various evasion techniques have also been observed to hide a contact deeply in the respective PIP. Equipped with such evasion tactics, it would become challenging for the OSN platform to extract the contacts from a detected PIP. Such contact-wise evasion techniques can be grouped into 3 categories, as detailed below.
Visual representation and steganography. The visual representation of contacts involves presenting contact information as visual elements in images or even videos. This can be further strengthened by the use of human-written text (Figure 4(a)), which makes it difficult for even OCR systems to recognize and extract the respective contact information. Embedding contact information within seemingly innocuous elements is also a way of stealthy promotion, e.g., embedding the contact in the poll options instead of the tweet main text. Additionally, illicit operators may also embed contact information within any frames of innocuous videos. Currently, our PIP contact extractor supports only the retrieval of contacts from text inputs, namely, the PIP text and the profile text of a PIP account. We leave it as our future work to extract contacts from media files and other seemingly innocuous components of a PIP.
Private messaging. By requesting potential buyers to send private messages for more information (e.g., Figure 8(a) in Appendix A), illicit operators can avoid the direct exposure of their contact information in OSNs. This tactic enables them to selectively respond to potential buyers and increase the effectiveness of their promotional efforts.
Contact preference across PIP categories. Regarding contact selection, illicit promotion campaigns of different categories exhibit distinct preferences. Table 10 presents, for each PIP category, how the extracted contact entities distribute across well-known contact types. Porn, money laundering, and weapon sales predominantly prefer Telegram accounts, which are associated with 43.60%, 68.17%, 73.36% of PIPs of the corresponding category, respectively. For harassment activities, QQ emerges as the most favored contact method, accounting for 56.23% of harassment PIPs. Besides, for data theft & leakage and fake documents, WeChat is much more popular than others.
Category1 Telegram Wechat QQ LINE WA2
Pornography 43.6% 30.24% 34.02% 2.49% 0.02%
Illegal Drug 36.99% 24.91% 13.79% 14.73% 1.12%
Gambling 42.45% 41.39% 24.73% 2.96% 0.97%
Surrogacy 22.92% 67.76% 14.26% 0.03% 1.38%
Harassment 25.60% 22.78% 56.23% 0.01% 0.00%
Money Laundering 68.17% 19.77% 17.91% 0.04% 0.24%
Weapon Sales 73.36% 18.80% 7.86% 0.1% 2.5%
Data Theft and Leakage 35.75% 58.91% 11.51% 0.02% 1.50%
FFD 2 16.64% 68.76% 14.81% 0.47% 0.00%
Crowdturfing 55.23% 44.18% 11.04% 0.00% 0.00%
  • 1
    Note that one PIP may include more than one contact.
  • 2
    WA denotes WhatsApp and FFD is short for Forgery and Fake Documents.
The adoption of novel IM platforms. In addition to the widely used communication platforms, we have also observed the adoption of newly emerged end-to-end encrypted communication platforms, such as Wickr(wic, 2023), BatChat(bat, 2023), and Potato Chat(pot, 2023). All these platforms support end-to-end encrypted communication, just like Telegram and WhatsApp. Furthermore, we find that these secure communication platforms support one or more novel security features which can facilitate more stealthy communication compared with aforementioned well-adopted ones.
Specifically, BatChat provides secret chat mode, in which, both parties’ avatars are encoded, making it difficult to identify the participants. Additionally, the platform disables the ability to take screenshots or forward chat messages, further safeguarding the content of the conversation. Besides, the secret chat mode of Potato Chat supports even more security features. For instance, Potato enforces message deletion by instructing the recipient’s app to delete messages when the sender removes them. Furthermore, users can set self-destruction timers for messages, photos, videos, and files, automatically removing the content from both devices after a specified time.
@@
Threat alerts from VirusTotal. To profile the threats of websites embedded in PIPs, we retrieved and analyzed their threat reports from VirusTotal (vir, 2023), a widely recognized open threat exchange platform. Due to the rate limit of VT, we only randomly sampled 20k PIP website URLs (corresponding to 7,391 FQDNs) for analysis. As shown in Table 11, 0.80% URLs have one or more alarms triggered, while it is 2.46% for FQDNs. Furthermore, among these alarmed URLs, 11.38% are considered as malware websites, while 24.39% are considered as phishing websites.
Table 11. Statistics of VirusTotal reports for PIP websites.
Category Count Reported Alarmed Malware Phishing
URLs 20,000 7.88% 0.80% 0.07% 0.15%
FQDNs 7,391 69.30% 2.46 % 0.32% 0.70%

### v3

@@ 4. Posts of Illicit Promotion
The ratio of PIPs to the Twitter stream. Another interesting question is to profile what fraction of general tweets are PIPs. As described in §3.1, we utilized daily tweet snapshots from the Twitter Archiving Project to evaluate the generalizability of our binary PIP classifier, which achieves a precision of 99.33%. We then selected the daily tweet snapshot of the 6th day of each month between January 2023 and June 2023, and applied our binary PIP classifier to these 6 snapshots. As presented in Table 3, the ratio of PIPs ranges from 2.69% to 4.75%, which, as a concerning fraction, highlights the prevalence of PIPs on Twitter.
Categories. Leveraging the multiclass PIP classifier (§3.2), all the captured PIPs have been predicted into one of well-defined 11 categories of illicit goods and services. The distribution of PIPs across these categories is listed in Table 4. And we can see the top 3 categories are Pornography,Gambling and Illegal Drug, which account for 91.17% PIPs in sum. While more observations are presented in Appendix F due to page limit, one thing that deserves attention is the discovery of a large amount of PIPs promoting products or services of child pornography, which have severely violated Twitter’s content policy regarding child sexual exploitation (chi, 2023). For instance, Figure4(a) in Appendix F is an example of child porn PIP posted in an evasive manner. The tweet text does not reveal itself as pornography, nor does it specify how to access the pornography resources. Instead, the hashtags and the attached image jointly promote child pornography as well as stating that the Telegram group can be found in the post author’s account profile (Figure 4(b)).
Table 4. The distribution of PIPs across illicit categories.
@@
Gambling 13.34% Forgery/Fake Documents 0.43%
@@
Data Theft/Leakage 2.03% Harassment 0.18%
@@
We further investigate the accounts that have all posts detected as PIPs, which comprise 7.87% of all PIP accounts. Our analysis reveals that such accounts tend to be dedicated to promoting a specific category of PIPs. Also, some accounts have posted very few tweets since registration, but rely primarily on their profiles rather than tweets for illicit promotion. Also, among top accounts with the largest volume of PIPs, we observe that a Twitter account can post hundreds of thousands of PIPs without getting banned by the platform. For instance, a PIP account registered in February, 2020 has posted over 198k PIPs in the category of data leakage, which strongly indicates the limitation of Twitter’s content moderation.
@@
PIP distribution across natural languages. To recognize the natural language of a PIP, fastText  [^2] https://fasttext.cc/, a library for text representation and classification from Facebook’s AI Research (FAIR) lab, is applied. As the result, most PIPs (over 99%) are grouped into 5 major natural languages including Chinese (48.75%), English (38.24%), Japanese (9.98%), Thai (1.70%) and Spanish (0.54%). In terms of illicit categories, the top 5 languages are similar to each other, as Pornography takes a major place in almost all of the top languages, followed by gambling or illegal drug. Considering products of these categories are illegal but can be of great demand in most countries, it’s reasonable to have observed such results.
Social engagement on PIPs. Over time, PIPs accumulate engagements such as likes, replies, retweets and quotes (i.e. retweets with comments). For our analysis on such social engagement on PIPs, please refer to Appendix I.
Finding I: Illicit promotion on Twitter is prevalent to a concerning extent, featuring a wide distribution across Twitter accounts, categories of illicit goods and services, and natural languages.
@@
Availability across online social networks To investigate availability of PIPs across other online social network, we conducted a cross-platform analysis for other three major platforms: YouTube, Facebook, and TikTok. For each PIP category, we randomly selected 20 PIP keywords which led to PIPs of the respective category identified on Twitter, and manually searched these platforms for PIPs if any. As shown in Table 5, PIPs are present on all these three platforms, albeit with varying degrees of prevalence. Particularly, Facebook has the highest hit ratio, with 83% of the sampled search keywords having at least one PIP located, and Youtube with the lowest at 35%. Besides, the availability of PIPs varies across platforms and categories. For example, the hit ratios of both gambling and data theft and leakage are higher on Tiktok than other subcategories. On the other hand, the promotion of weapon sales has a low hit ratio on all three platforms, with less than 30% of the sampled seeds having one or more PIPs observed. Interestingly, we found that some PIPs on different platforms share the same contact, indicating that they belong to the same promotion campaign.
Finding II: Illicit promotion is extensively observed across online social networks, highlighting itself as a cross-platform challenge.
@@ 5. Content Moderation Against Illicit Promotion
All OSNs under our study claim to have enforced strict content moderation (tik, 2023; twi, 2023b; you, 2023), in which case, violative posts should either be blocked from publishing or get unpublished once detected in a later time. However, The existence of so many violative PIPs of diverse illicit goods and services categories suggest respective OSNs especially Twitter fail to prevent PIPs from being posted and becoming visible to OSN users. We thus take a closer look into how PIPs evade the content moderation of Twitter, as detailed below.
Evasion tactics of PIPs. To comprehensively analyze the evasion tactics of PIPs, we randomly sample 500 PIPs for each illicit category, manually look into each of them to distll tactics, and implement quantitative analysis on the whole PIP dataset to verify the applicability. The evasion tactics adopted by miscreants can be summarized into three categories: I) Leveraging benign and popular hashtags; II) Using jargon words; III) Embedding illicit promotion messages into any component of a PIP but not its main text. Detailed explanation and examples can be found in Appendix H.
Finding III: Illicit promotion campaigns have adopted various evasion tactics, likely in an attempt to evade content moderation of the Twitter platform.
Table 6. The evasion rate of PIPs across time.
@@
Content moderation of Twitter. To further profile the effectiveness of Twitter’s content moderation towards published PIPs, we carried out revisiting for captured PIPs. Specifically, 50,000 PIPs hunted in each round are sampled, and their availability are tested by revisiting them periodically (usually at a weekly pace). Given sampled $PIP_{date_{p}}$ which are first posted on date $date_{p}$, their evasion rate $ER_{date_{p}\text{, }date_{r}}$ is defined as the ratio of PIPs that are still reachable when being revisited on date $date_{r}$. Table 6 presents the revisiting results between April 10 and May 1 in 2023, for PIPs posted between Oct 24, 2022 and March 12, 2023. And we can see PIPs posted between Oct 24-30, 2022 have 21.46% still reachable when revisiting 6 month later in the end of April 2023. Also, more than 90% PIPs can survive the first two months since being published.
@@
We then investigate why some PIPs become unavailable during revisits and find out that most are due to account suspension. Specifically, Twitter returns one of six error messages if the PIP is unavailable. These messages can tell us the underlying reasons of PIP unavailability. For instance, a PIP may be unpublished due to the suspension of the parental account, in which case, the error message will be ”This Tweet is from a suspended account.”. Other reasons include page non-existence, deletion by the author, account non-existence and violation of Twitter’s rules. Among unreachable PIPs during revisits, 91.59% are due to account suspension, and 6.22% are due to page non-existence. In summary, we can see that Twitter works in a continuous manner to detect and suspend PIP accounts, resulting in tweets (including PIPs) of the detected PIP accounts also become unpublished.
@@
However, as many PIPs can survive for a long period, we further investigate the difference between unpublished PIPs and surviving ones. To answer this question, PIPs are sampled and divided into two groups depending on the length of their evasion time as observed during revisiting. We then compare both groups with regards to various aspects, e.g., PIP categories, the text syntactic and semantics, the writing style, characteristics of the posting accounts. The only significant difference we have observed resides in the number of PIPs posted by the parental Twitter accounts. For the group of PIPs with a short lifetime, their parental Twitter accounts have 379 PIPs observed on average in our dataset. On the contrary, it is only 62 for the group of PIPs with much longer evasion time. A reasonable explanation is that the more PIPs a Twitter account has posted, the more likely it will be captured and thus suspended by the platform, in which case, all its tweets will also be unavailable, including the PIPs. To further profile this observation, we sampled 153,921 out of all the observed 580,530 PIP accounts, and revisited them during April 15-19, 2023. By then, only 78.95% accounts were still available while all the others got suspended. Comparing alive PIP accounts with suspended ones reveals that alive PIP accounts have 3 PIPs observed on average while blocked ones have 87.
Finding IV: An arms race is observed between illicit promotion campaigns and Twitter’s continuous content moderation, as the result of which, almost 80% PIPs have got banned six months later after being published while on the other hand, 90% PIPs can survive the first two months.
@@ 6.1. PIP Contacts
The adoption of novel IM platforms. In addition to the widely used communication platforms, we have also observed the adoption of newly emerged end-to-end encrypted communication platforms, such as Wickr(wic, 2023), BatChat(bat, 2023), and Potato Chat(pot, 2023). All these platforms support end-to-end encrypted communication, just like Telegram and WhatsApp. Furthermore, we find that these secure communication platforms support one or more novel security features which can facilitate more stealthy communication compared with aforementioned well-adopted ones. Specifically, BatChat provides secret chat mode, in which, both parties’ avatars are encoded, making it difficult to identify the participants. Additionally, the platform disables the ability to take screenshots or forward chat messages, further safeguarding the content of the conversation. Besides, the secret chat mode of Potato Chat supports even more security features. For instance, Potato enforces message deletion by instructing the recipient’s app to delete messages when the sender removes them. Furthermore, users can set self-destruction timers for messages, photos, videos, and files, automatically removing the content from both devices after a specified time.
@@
For analysis concerning contact preference across PIP categories and threat alerts from VirusTotal, please refer to Appendix J.